# SAVA: Scalable Learning-Agnostic Data Valuation

### v1

@@ Abstract
Selecting suitable data for training machine learning models is crucial since large, web-scraped, real datasets contain noisy artifacts that affect the quality and relevance of individual data points. These artifacts will impact the performance and generalization of the model. We formulate this problem as a data valuation task, assigning a value to data points in the training set according to how similar or dissimilar they are to a clean and curated validation set. Recently, LAVA  (Just et al.,, 2023) successfully demonstrated the use of optimal transport (OT) between a large noisy training dataset and a clean validation set, to value training data efficiently, without the dependency on model performance. However, the LAVA algorithm requires the whole dataset as an input, this limits its application to large datasets. Inspired by the scalability of stochastic (gradient) approaches which carry out computations on batches of data points instead of the entire dataset, we analogously propose SAVA, a scalable variant of LAVA with its computation on batches of data points. Intuitively, SAVA follows the same scheme as LAVA which leverages the hierarchically defined OT for data valuation. However, while LAVA processes the whole dataset, SAVA divides the dataset into batches of data points, and carries out the OT problem computation on those batches. We perform extensive experiments, to demonstrate that SAVA can scale to large datasets with millions of data points and doesn’t trade off data valuation performance.
@@ 1 Introduction
Neural scaling laws empirically show that the generalization error decreases according to a power law as the data a model trains on increases. This has been shown for natural language processing, vision, and speech (Kaplan et al.,, 2020; Henighan et al.,, 2020; Rosenfeld et al.,, 2019; Zhai et al.,, 2022; Radford et al.,, 2023). However, training neural networks on larger and larger datasets for moderate improvements in model accuracy is inefficient. Furthermore, production neural network models need to be continuously updated given new utterances that enter into common everyday parlance (Lazaridou et al.,, 2021; Baby et al.,, 2022). It has been shown both in theory and in practice that sub-power law, exponential scaling of model performance with dataset size is possible by carefully selecting informative data and pruning uninformative data (Sorscher et al.,, 2022), and that generalization improves with training speed (Lyle et al.,, 2020). Therefore, valuing and selecting data points that are informative: which have not been seen by the model, which do not have noisy labels, and which are relevant to the task we want to solve—are not outliers—can help to not only decrease training times, and reduce compute costs, but also improve overall test performance (Mindermann et al.,, 2022; Tirumala et al.,, 2023).
@@
Figure 1: Overview of the proposed SAVA method. Instead of solving a single, but expensive, OT problem on the whole dataset, SAVA considers solving multiple cheaper smaller OT problems on batches of data points, and integrates the resulting calibrated gradients together for a data valuation task. $\text{OT}(\bar{\mu_{t}},\bar{\mu_{v}})$ denotes the OT cost and $\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})$ is its associated optimal transport plan. $\text{OT}(\mu_{B_{i}},\mu_{B_{j}^{\prime}})$ is the OT between a uniform distribution of points in the batch $B_{i}$ from the training set and the batch $B^{\prime}_{j}$ in the validation set. $s_{k}$ is SAVA’s final valuation score for training data point $z_{k}$. We provide a visualization of these artifacts generated by SAVA in Figure 12.
@@
Popular data selection and data pruning methods use variations of the model loss to value data points (Jiang et al.,, 2019; Pruthi et al.,, 2020; Paul et al.,, 2021). Crucially, these methods depend on the model used, and they are vulnerable to prioritizing data points with noisy labels or noisy features; data points that do not resemble the target validation set. When treating data valuation as a function of model performance, we introduce a dependency on a neural network model. This is the case when valuing a point using the leave-one-out (LOO) error, i.e., the change of model performance when the point is omitted from training. To rid our dependence on a neural network model, a promising idea is to leverage optimal transport (OT) between a training distribution and a clean validation distribution as a proxy to directly measure the value of data points in the training set in a model-agnostic fashion. In particular, the validation performance of each point in the training set can be estimated using the hierarchically defined Wasserstein between the training and the validation set (Alvarez-Melis and Fusi,, 2020; Just et al.,, 2023). LAVA (Just et al.,, 2023) has been shown to successfully value data by measuring the sensitivity of the hierarchically defined Wasserstein between training data points and validation distributions in a model-agnostic fashion. However, LAVA requires significant RAM consumption since its memory complexity grows quadratically $\mathcal{O}(N^{2})$ with the dataset size $N$. This hinders LAVA from scaling to large datasets.
@@
In this paper, we present SAVA, a scalable variant of LAVA, for data valuation. Our method completely addresses the bottleneck of RAM requirements in LAVA. Intuitively, SAVA performs the OT computations on batches instead of on the entire dataset like LAVA (hence the analogy to the stochastic approach to gradient computation). Specifically, SAVA uses ideas from hierarchical optimal transport (Yurochkin et al.,, 2019; Lee et al.,, 2019) to enable OT calculations on batches instead of the entire dataset. We can scale up OT-based data valuation using SAVA to large real-world web-scrapped datasets. On benchmark problems SAVA performs comparably to LAVA while being able to scale to datasets two orders of magnitude larger without memory issues, while LAVA is limited due to hardware memory constraints.
@@
1. We introduce a novel scalable data valuation method called SAVA that leverages hierarchical optimal transport (Yurochkin et al.,, 2019; Lee et al.,, 2019) which performs OT computations on batches of data points, enabling OT-based data valuation to large datasets.
@@ 2.1 Optimal Transport for Labeled Datasets
Let ${\mathcal{X}}$ be the feature space, and $V$ be the number of labels. We write $f_{t}:{\mathcal{X}}\mapsto\left\{0,1\right\}^{V}$ and $f_{v}:{\mathcal{X}}\mapsto\left\{0,1\right\}^{V}$ for the labeling functions for training and validation data respectively. Given the training set ${\mathbb{D}}_{t}=\left\{(x_{i},f_{t}(x_{i}))\right\}_{i=1}^{N}$ and the validation set ${\mathbb{D}}_{v}=\left\{(x^{\prime}_{i},f_{v}(x^{\prime}_{i}))\right\}_{i=1}^{N^{\prime}}$, the corresponding measures for sets ${\mathbb{D}}_{t},{\mathbb{D}}_{v}$ are $\mu_{t}(x,y)=\frac{1}{N}\sum_{i=1}^{N}\delta_{(x_{i},y_{i})}$ and $\mu_{v}(x^{\prime},y^{\prime})=\frac{1}{N^{\prime}}\sum_{i=1}^{N^{\prime}}\delta_{(x^{\prime}_{i},y^{\prime}_{i})}$ respectively where $\delta$ is the Dirac function, and $y,y^{\prime}$ are labels of $x,x^{\prime}$ respectively. For simplicity, let ${\mathcal{Z}}=({\mathcal{X}},{\mathcal{Y}})$ where ${\mathcal{Y}}$ is the space of labels. For ease of reading, we summarize all notations in Appendix Table 1.
Following Alvarez-Melis and Fusi, (2020), we compute the distance between two labels by leveraging the OT distance between the conditional distributions of the features given each label. More concretely, for label $y_{t}$ in $\mu_{t}$, the conditional distribution of the features given $y_{t}$ in $\mu_{t}$ is as $\mu_{t}(\cdot|y_{t})=\frac{\mu_{t}(\cdot)I[f_{t}(\cdot)=y_{t}]}{\int\mu_{t}(\cdot)I[f_{t}(\cdot)=y_{t}]}$, where $I$ is the indicator function (similarly for label $y_{v}$ in $\mu_{v}$). Let d be the metric of the feature space ${\mathcal{X}}$, e.g., we use the Euclidean distance for d in practice. The distance between labels $y_{t}$ and $y_{v}$ is $\textnormal{OT}_{{\textnormal{d}}}(\mu_{t}(\cdot|y_{t}),\mu_{v}(\cdot|y_{v}))$, i.e., the metric of the label space ${\mathcal{Y}}$. Thus, we can compute the transport cost function between feature-label pairs in ${\mathcal{Z}}=({\mathcal{X}},{\mathcal{Y}})$ as follows
${\textnormal{C}}((x_{t},y_{t}),(x_{v},y_{v}))={\textnormal{d}}(x_{t},x_{v})+c\,\textnormal{OT}_{{\textnormal{d}}}(\mu_{t}(\cdot|y_{t}),\mu_{v}(\cdot|y_{v})),$ (1)
where $c>0$ is a weight coefficient. Given the cost matrix C for labeled data points (i.e., feature-label points in ${\mathcal{Z}}$) in Eq. (1), we can use the OT on the represented measures $\mu_{t},\mu_{v}$ to compute the distance between the training and validation sets, i.e., $d({\mathbb{D}}_{t},{\mathbb{D}}_{v})$, without relying on external models or parameters. It is worth noting that the cost matrix is hierarchically defined; it is dependent on the solution of the OT problem between labels Eq. (1). More concretely, we have
@@
where $\Pi(\mu_{t},\mu_{v})$ is the set of transportation couplings which have marginals as $\mu_{t}$ and $\mu_{v}$. To simplify notations, we drop C, and use OT when the context is clear. We further write $\pi^{*}$ for the optimal transport plan in Eq. (2). To deal with high computational complexity, i.e., super cubic w.r.t. the number of supports of input measures, an efficient approach utilizes entropic regularization (Cuturi,, 2013) to reduce its complexity into quadratic, defined as
@@ 2.2 LAVA: Data valuation via calibrated OT gradients
As mentioned in Just et al., (2023), the OT distance is known to be insensitive to small differences while also being not robust to large deviations. This feature is naturally suitable for detecting abnormal data points, i.e., disregarding normal variations in distances between clean data while being sensitive to abnormal distances of outlying points. Therefore, the gradient of the OT distance w.r.t. the probability mass associated with each point can be leveraged as a surrogate to measure the contribution of that point. More precisely, the gradient of the OT distance w.r.t. the probability mass of data points in the two datasets can be expressed:
@@
Note that we can leverage the Sinkhorn algorithm to solve the entropic regularized OT, i.e., $\textnormal{OT}_{\varepsilon}(\mu_{t},\mu_{v})$, and obtain the optimal dual variables $(f^{*},g^{*})$ as its by-product, for computing the gradient of entropic regularized OT w.r.t. the mass of support data points of input measures.
The dual solution is unique, up to a constant due to the redundant constraint $\sum_{i=1}^{N}\mu_{t}(z_{i})=\sum_{i=1}^{M}\mu_{v}(z^{\prime}_{i})=1$. Therefore, for measuring the gradients of the OT w.r.t. the probability mass of a given data point in each dataset, Just et al., (2023) calculate the calibrated gradients (i.e., a sum of all elements equals to a zero vector) as
$\frac{\partial\textnormal{OT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{i})}=f^{*}_{i}-\sum_{j\in\left\{1,...,N\right\}\textbackslash i}\frac{f^{*}_{j}}{N-1},$ (6)
The calibrated gradients predict how the OT distance changes as more probability mass is shifted to a given data point. This can be interpreted as a measurement of the contribution of the data point to the OT. Additionally, if we want a training set to match the distribution of the validation dataset, then either removing or reducing the mass of data points with large positive gradients, while increasing the mass of data points with large negative gradients can be expected to reduce their OT distance. Therefore, as demonstrated in Just et al., (2023), the calibrated gradients provide a powerful tool to detect and prune abnormal or irrelevant data in various applications.
Memory Limitation. While being used with the current best practice, the Sinkhorn algorithm for entropic regularized OT (Cuturi,, 2013) still runs in quadratic memory complexity $O(N^{2})$ with the dataset size $N$, as it requires performing operations on the entire dataset, using the full pairwise cost matrix. Consequently, the memory and RAM requirements for the Sinkhorn algorithm primarily depend on the dataset size $N$. Additionally, notice that a dense square (float) matrix of size $N=10^{5}$ will require at least $74$ GB of RAM and $N=10^{6}$ will take $7450$ GB of RAM which is prohibitively expensive. Therefore, LAVA is limited to small datasets.
@@
It is worth noting that the OT-based data valuation we introduced in this section focuses on the gradient of the OT instead of the OT distance itself. Therefore, some popular scalable OT approaches, e.g., sliced-Wasserstein (Rabin et al.,, 2011), tree-sliced-Wasserstein (Le et al.,, 2019), or Sobolev transport (Le et al.,, 2022), may not be suitable to data valuation. In the next section, we introduce our novel scalable approach for computing the gradient of the OT using hierarchical OT approach (Yurochkin et al.,, 2019; Lee et al.,, 2019). We focus on the problem of data valuation but our work can be readily applied for other large-scale applications where the gradient of the OT is required.
@@ 3 SAVA: Scalable Data Valuation
We present SAVA, a scalable data valuation method, scaling LAVA to a large-scale dataset. Instead of solving a single (but expensive) OT problem for distributions on the entire datasets, i.e., $\textnormal{OT}(\mu_{t},\mu_{v})$ in LAVA with the pairwise cost matrix size ${\mathbb{R}}^{N\times N^{\prime}}$, we consider solving multiple (cheaper) OT problems for distributions on batches of data points. For this purpose, our algorithm performs data valuation on two levels of hierarchy: across batches, and across data points within two batches. Thus, SAVA can complement LAVA for large-scale applications.
@@
Input: a threshold $\tau$, $\epsilon$ for Sinkhorn algorithm, let $z=(x,y)$
@@
6            543 Compute ${\textnormal{C}}_{kl}\bigl{(}B_{i},B^{\prime}_{j}\bigr{)},\forall k\in[N_{i}],\forall l\in[N^{\prime}_{j}]$ by using Eq. (8). Compute ${\color[rgb]{1,0,0}f^{*}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})},g^{*}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})$ by solving $\textnormal{OT}_{{\textnormal{C}}}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})$. Set ${\color[rgb]{.5,0,.5}\bar{{\textnormal{C}}}_{ij}(\bar{\mu}_{t},\bar{\mu}_{v})}=\textnormal{OT}_{{\textnormal{C}}}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})$.           // i.e., distance $d({B}_{i},{B}^{{}^{\prime}}_{j})$ on batches.
@@
15            1413 Compute $\frac{\partial\textnormal{OT}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})}{\partial\mu_{{B}_{i}}(z_{k})},\,\,\forall j\in[K_{v}]$ using Eq. (14). Compute $s_{k}=\frac{\partial\textnormal{HOT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{l})}$ using Eq. (13). // valuation score for data point $z_{k}\in B_{i}$.
@@
Hierarchical OT. We follow the idea in hierarchical OT approach (Yurochkin et al.,, 2019; Lee et al.,, 2019) to partition the training dataset ${\mathbb{D}}_{t}$ of $N$ samples into $K_{t}$ disjoint batches ${B}=\left\{{B}_{i}\right\}_{i=1}^{K_{t}}$. Similarly, for the validation set ${\mathbb{D}}_{v}$ of $N^{\prime}$ samples into $K_{v}$ disjoint batches ${B}^{\prime}=\left\{{B}^{\prime}_{j}\right\}_{j=1}^{K_{v}}$. Additionally, for all $i\in[K_{t}],j\in[K_{v}]$, let the number of samples in batches $B_{i},B^{\prime}_{j}$ as $N_{i},N^{\prime}_{j}$ respectively. The corresponding measures of the training and validation sets w.r.t. the batches are defined as: $\bar{\mu}_{t}({B})=\frac{1}{K_{t}}\sum_{i=1}^{K_{t}}\delta_{({B}_{i})}$ and $\bar{\mu}_{v}({B}^{\prime})=\frac{1}{K_{v}}\sum_{j=1}^{K_{v}}\delta_{({B}^{\prime}_{j})}$ respectively. We then define a distance between the datasets as the hierarchical optimal transport (HOT) between the measures $\mu_{t},\mu_{v}$ as OT distance for corresponding represented measures on batches, i.e., $\bar{\mu}_{t}$ and $\bar{\mu}_{v}$ as in Sec. 2.1 as follows:
@@
Our goal is to estimate the gradient $\frac{\partial d(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{k})}=\frac{\partial\textnormal{HOT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{k})}$ where $\textnormal{HOT}(\mu_{t},\mu_{v})$ is defined in the Eq. (7). Computing this derivative involves two OT estimation steps including (i) OT between individual data points within two batches to compute $d({B}_{i},{B}^{\prime}_{j}):=\textnormal{OT}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})$, where $\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}}$ are corresponding measures for batches ${B}_{i},{B}^{\prime}_{j}$ respectively, and subsequently a cost matrix ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}$ on pairwise batches for input measures over batches; (ii) $d(\mu_{t},\mu_{v})=\textnormal{HOT}(\mu_{t},\mu_{v}):=\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}}(\bar{\mu}_{t},\bar{\mu}_{v})$.

### v2

@@ Abstract
Selecting data for training machine learning models is crucial since large, web-scraped, real datasets contain noisy artifacts that affect the quality and relevance of individual data points. These noisy artifacts will impact model performance. We formulate this problem as a data valuation task, assigning a value to data points in the training set according to how similar or dissimilar they are to a clean and curated validation set. Recently, LAVA  (just2023lava) demonstrated the use of optimal transport (OT) between a large noisy training dataset and a clean validation set, to value training data efficiently, without the dependency on model performance. However, the LAVA algorithm requires the entire dataset as an input, this limits its application to larger datasets. Inspired by the scalability of stochastic (gradient) approaches which carry out computations on batches of data points instead of the entire dataset, we analogously propose SAVA, a scalable variant of LAVA with its computation on batches of data points. Intuitively, SAVA follows the same scheme as LAVA which leverages the hierarchically defined OT for data valuation. However, while LAVA processes the whole dataset, SAVA divides the dataset into batches of data points, and carries out the OT problem computation on those batches. We perform extensive experiments, to demonstrate that SAVA can scale to large datasets with millions of data points and doesn’t trade off data valuation performance.
@@ 1 Introduction
Neural scaling laws empirically show that the generalization error decreases according to a power law as the data a model trains on increases. This has been shown for natural language processing, vision, and speech (kaplan2020scaling; henighan2020scaling; rosenfeld2019constructive; zhai2022scaling; radford2023robust). However, training neural networks on larger and larger datasets for moderate improvements in model accuracy is inefficient. Furthermore, production neural network models need to be continuously updated given new utterances that enter into common everyday parlance (lazaridou2021mind; baby2022incremental). It has been shown both in theory and in practice that sub-power law, exponential scaling of model performance with dataset size is possible by carefully selecting informative data and pruning uninformative data (sorscher2022beyond), and that generalization improves with training speed (lyle2020bayesian). Therefore, valuing and selecting data points that are informative: which have not been seen by the model, which do not have noisy labels, and which are relevant to the task we want to solve—are not outliers—can help to not only decrease training times, and reduce compute costs, but also improve overall test performance (mindermann2022prioritized; tirumala2023d4).
@@
Figure 1: Overview of the proposed SAVA method. Instead of solving a single, but expensive, OT problem on the whole dataset, SAVA considers solving multiple cheaper smaller OT problems on batches of data points, and integrates the resulting calibrated gradients together for a data valuation task. $\text{OT}(\bar{\mu_{t}},\bar{\mu_{v}})$ denotes the OT cost and $\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})$ is its associated optimal transport plan. $\text{OT}(\mu_{B_{i}},\mu_{B_{j}^{\prime}})$ is the OT between a uniform distribution of points in the batch $B_{i}$ from the training set and the batch $B^{\prime}_{j}$ in the validation set. $s_{k}$ is SAVA’s final valuation score for training data point $z_{k}$. We provide a visualization of these artifacts generated by SAVA in LABEL:fig:sava_alg_artifacts.
@@
Popular data selection and data pruning methods use variations of the model loss to value data points (jiang2019accelerating; pruthi2020estimating; paul2021deep). Crucially, these methods depend on the model used, and they are vulnerable to prioritizing data points with noisy labels or noisy features; data points that do not resemble the target validation set. When treating data valuation as a function of model performance, we introduce a dependency on a neural network model. This is the case when valuing a point using the leave-one-out (LOO) error, i.e., the change of model performance when the point is omitted from training. To rid our dependence on a neural network model, a promising idea is to leverage optimal transport (OT) between a training distribution and a clean validation distribution as a proxy to directly measure the value of data points in the training set in a model-agnostic fashion. In particular, the validation performance of each point in the training set can be estimated using the hierarchically defined Wasserstein between the training and the validation set (alvarez2020geometric; just2023lava). LAVA (just2023lava) has been shown to successfully value data by measuring the sensitivity of the hierarchically defined Wasserstein between training data points and validation distributions in a model-agnostic fashion. However, LAVA requires significant RAM consumption since its memory complexity grows quadratically $\mathcal{O}(N^{2})$ with the dataset size $N$. This hinders LAVA from scaling to large datasets.
@@
In this paper, we present SAVA, a scalable variant of LAVA, for data valuation. Our method completely addresses the bottleneck of RAM requirements in LAVA. Intuitively, SAVA performs the OT computations on batches instead of on the entire dataset like LAVA (hence the analogy to the stochastic approach to (sub)gradient computation). Specifically, SAVA uses ideas from hierarchical optimal transport (yurochkin2019hierarchical; lee2019hierarchical) to enable OT calculations on batches instead of the entire dataset. We can scale up OT-based data valuation using SAVA to large real-world web-scrapped datasets. On benchmark problems SAVA performs comparably to LAVA while being able to scale to datasets two orders of magnitude larger without memory issues, while LAVA is limited due to hardware memory constraints.
@@
1. We introduce a novel scalable data valuation method called SAVA that leverages the (sub)gradient of hierarchical optimal transport (yurochkin2019hierarchical; lee2019hierarchical) which performs OT computations on batches of data points, enabling OT-based data valuation to large datasets.
2. We correct the essential theoretical result on the trade-off for using entropic regularization to compute the OT (sub)gradient in LAVA (just2023lava, Theorem 2).111see Appendix LABEL:appendix_issue_theorem2_lava for the discussion Consequently, by building upon our refined theory, we derive the exact trade-off to estimate the (sub)gradient of hierarchical OT with entropic regularization in our framework.
@@ 2.1 Optimal Transport for Labeled Datasets
Let ${\mathcal{X}}$ be the feature space, and $V$ be the number of labels. We write $f_{t}:{\mathcal{X}}\mapsto\left\{0,1\right\}^{V}$ and $f_{v}:{\mathcal{X}}\mapsto\left\{0,1\right\}^{V}$ for the labeling functions for training and validation data respectively. Given the training set ${\mathbb{D}}_{t}=\left\{(x_{i},f_{t}(x_{i}))\right\}_{i=1}^{N}$ and the validation set ${\mathbb{D}}_{v}=\left\{(x^{\prime}_{i},f_{v}(x^{\prime}_{i}))\right\}_{i=1}^{N^{\prime}}$, the corresponding measures for sets ${\mathbb{D}}_{t},{\mathbb{D}}_{v}$ are $\mu_{t}(x,y)=\frac{1}{N}\sum_{i=1}^{N}\delta_{(x_{i},y_{i})}$ and $\mu_{v}(x^{\prime},y^{\prime})=\frac{1}{N^{\prime}}\sum_{i=1}^{N^{\prime}}\delta_{(x^{\prime}_{i},y^{\prime}_{i})}$ respectively where $\delta$ is the Dirac function, and $y,y^{\prime}$ are labels of $x,x^{\prime}$ respectively. For simplicity, let ${\mathcal{Z}}=({\mathcal{X}},{\mathcal{Y}})$ where ${\mathcal{Y}}$ is the space of labels. For ease of reading, we summarize all notations in Appendix Table LABEL:table:notation.
Following alvarez2020geometric, we compute the distance between two labels by leveraging the OT distance between the conditional distributions of the features given each label, i.e., $\mu_{t}(\cdot|y_{t})=\frac{\mu_{t}(\cdot)I[f_{t}(\cdot)=y_{t}]}{\int\mu_{t}(\cdot)I[f_{t}(\cdot)=y_{t}]}$ for label $y_{t}$ in $\mu_{t}$, where $I$ is the indicator function.
Let d be the metric of the feature space ${\mathcal{X}}$, e.g., the Euclidean distance. The distance between labels $y_{t}$ and $y_{v}$ is $\textnormal{OT}_{{\textnormal{d}}}(\mu_{t}(\cdot|y_{t}),\mu_{v}(\cdot|y_{v}))$, i.e., the metric of the label space ${\mathcal{Y}}$. Consequently, the cost between feature-label pairs in ${\mathcal{Z}}=({\mathcal{X}},{\mathcal{Y}})$ is ${\textnormal{C}}((x_{t},y_{t}),(x_{v},y_{v}))={\textnormal{d}}(x_{t},x_{v})+c\,\textnormal{OT}_{{\textnormal{d}}}(\mu_{t}(\cdot|y_{t}),\mu_{v}(\cdot|y_{v}))$, where $c>0$ is a weight coefficient. Therefore, with the cost matrix C, we can use the OT on the represented measures $\mu_{t},\mu_{v}$ to compute the distance between the training and validation sets, i.e., $d({\mathbb{D}}_{t},{\mathbb{D}}_{v})$, without relying on external models or parameters as follows
@@
where $\Pi(\mu_{t},\mu_{v})$ is the set of transportation couplings with marginals as $\mu_{t}$ and $\mu_{v}$. To simplify notations, we drop C, and use OT when the context is clear. We further write $\pi^{*}$ for the optimal transport plan in Eq. (1). In practice, we leverage the entropic regularization (Cuturi-2013-Sinkhorn) to reduce the OT complexity into quadratic (from super cubic) w.r.t. the number of input supports, defined as
@@ 2.2 LAVA: Data valuation via calibrated OT gradients
As mentioned in just2023lava, the OT distance is known to be insensitive to small differences while also being not robust to large deviations. This feature is naturally suitable for detecting abnormal data points, i.e., disregarding normal variations in distances between clean data while being sensitive to abnormal distances of outlying points. Therefore, the (sub)gradient of the OT distance w.r.t. the probability mass associated with each point can be leveraged as a surrogate to measure the contribution of that point. More precisely, the (sub)gradient of the OT distance w.r.t. the probability mass of data points in the two datasets can be expressed:
@@
The dual solution is unique, up to a constant due to the redundant constraint $\sum_{i=1}^{N}\mu_{t}(z_{i})=\sum_{i=1}^{M}\mu_{v}(z^{\prime}_{i})=1$. Therefore, for measuring the subgradients of the OT w.r.t. the probability mass of a given data point in each dataset, just2023lava we calculate the calibrated gradients (i.e., a sum of all elements equals to a zero vector) [^2] To remove the degree of freedom which comes from the fact that one among all row/column sum constraints for the transport polytope is redundant, among the OT subgradients, we fix the zero-sum subgradient following the convention in Cuturi-2014-Fast and the calibrated approach in LAVA (just2023lava). as
$\frac{\partial\textnormal{OT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{i})}=f^{*}_{i}-\sum_{j\in\left\{1,...,N\right\}\textbackslash i}\frac{f^{*}_{j}}{N-1}.$ (5)
The calibrated gradients predict how the OT distance changes as more probability mass is shifted to a given data point. This can be interpreted as a measurement of the contribution of the data point to the OT. Additionally, if we want a training set to match the distribution of the validation dataset, then either removing or reducing the mass of data points with large positive gradients, while increasing the mass of data points with large negative gradients can be expected to reduce their OT distance. Therefore, as demonstrated in just2023lava, the calibrated gradients provide a powerful tool to detect and prune abnormal or irrelevant data in various applications.
Memory Limitation. While being used with the current best practice, the Sinkhorn algorithm for entropic regularized OT (Cuturi-2013-Sinkhorn) still runs in quadratic memory complexity $O(N^{2})$ with the dataset size $N$, as it requires performing operations on the entire dataset, using the full pairwise cost matrix. Consequently, the memory and RAM requirements for the Sinkhorn algorithm primarily depend on the dataset size $N$. Additionally, notice that a dense square (float) matrix of size $N=10^{5}$ will require at least $74$ GB of RAM and $N=10^{6}$ will take $7450$ GB of RAM which is prohibitively expensive. Therefore, LAVA is limited to small datasets.
@@
The scalable OT-based data valuation we introduced in this section focuses on the (sub)gradient of the OT instead of the OT distance itself. Therefore, popular scalable OT approaches, sliced-Wasserstein (rabin2011wasserstein) or Sobolev transport (le2022st), may not be suitable. In the next section, we introduce our novel scalable approach for computing the (sub)gradient of the OT using hierarchical OT (yurochkin2019hierarchical; lee2019hierarchical). We focus on the problem of data valuation but our work can be applied to other large dataset applications where the (sub)gradient of the OT is required.
@@ 3 SAVA: Scalable Data Valuation
We present SAVA, a scalable data valuation method, scaling LAVA to large-scale datasets. Instead of solving a single (but expensive) OT problem for distributions on the entire datasets, i.e., $\textnormal{OT}(\mu_{t},\mu_{v})$ in LAVA with the pairwise cost matrix size ${\mathbb{R}}^{N\times N^{\prime}}$, we consider solving multiple (cheaper) OT problems for distributions on batches of data points. For this purpose, our algorithm performs data valuation on two levels of hierarchy: across batches, and across data points within two batches. Thus, SAVA can complement LAVA for large-scale applications.
@@
Input: a threshold $\tau$, $\varepsilon$ for Sinkhorn algorithm, let $z=(x,y)$
@@
6            543 Compute ${\textnormal{C}}_{kl}\bigl{(}B_{i},B^{\prime}_{j}\bigr{)},\forall k\in[N_{i}],\forall l\in[N^{\prime}_{j}]$ by using Eq. (7). Compute ${\color[rgb]{1,0,0}f^{*}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})},g^{*}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})$ by solving $\textnormal{OT}_{{\textnormal{C}}}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})$. Set ${\color[rgb]{.5,0,.5}\bar{{\textnormal{C}}}_{ij}(\bar{\mu}_{t},\bar{\mu}_{v})}=\textnormal{OT}_{{\textnormal{C}}}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})$.           // distance $d({B}_{i},{B}^{{}^{\prime}}_{j})$ on batches.
@@
15            1413 Compute $\frac{\partial\textnormal{OT}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})}{\partial\mu_{{B}_{i}}(z_{k})},\,\,\forall j\in[K_{v}]$ using Eq. (13). Compute $s_{k}=\frac{\partial\textnormal{HOT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{l})}$ using Eq. (12). // valuation score for $z_{k}\in B_{i}$.
@@
Hierarchical OT. We follow the idea in hierarchical OT approach (yurochkin2019hierarchical; lee2019hierarchical) to partition the training dataset ${\mathbb{D}}_{t}$ of $N$ samples into $K_{t}$ disjoint batches ${B}=\left\{{B}_{i}\right\}_{i=1}^{K_{t}}$. Similarly, for the validation set ${\mathbb{D}}_{v}$ of $N^{\prime}$ samples into $K_{v}$ disjoint batches ${B}^{\prime}=\left\{{B}^{\prime}_{j}\right\}_{j=1}^{K_{v}}$. Additionally, for all $i\in[K_{t}],j\in[K_{v}]$, let the number of samples in batches $B_{i},B^{\prime}_{j}$ as $N_{i},N^{\prime}_{j}$ respectively. The corresponding measures of the training and validation sets w.r.t. the batches are defined as: $\bar{\mu}_{t}({B})=\frac{1}{K_{t}}\sum_{i=1}^{K_{t}}\delta_{({B}_{i})}$ and $\bar{\mu}_{v}({B}^{\prime})=\frac{1}{K_{v}}\sum_{j=1}^{K_{v}}\delta_{({B}^{\prime}_{j})}$ respectively. We then define a distance between the datasets as the hierarchical optimal transport (HOT) between the measures $\mu_{t},\mu_{v}$ as OT distance for corresponding represented measures on batches, i.e., $\bar{\mu}_{t}$ and $\bar{\mu}_{v}$ as in Sec. 2.1 as follows:
@@
Our goal is to estimate the (sub)gradient $\frac{\partial d(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{k})}=\frac{\partial\textnormal{HOT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{k})}$ where $\textnormal{HOT}(\mu_{t},\mu_{v})$ is defined in the Eq. (6). Computing this derivative involves two OT estimation steps including (i) OT between individual data points within two batches to compute $d({B}_{i},{B}^{\prime}_{j}):=\textnormal{OT}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})$, where $\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}}$ are corresponding measures for batches ${B}_{i},{B}^{\prime}_{j}$ respectively, and subsequently a cost matrix ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}$ on pairwise batches for input measures over batches; (ii) $d(\mu_{t},\mu_{v})=\textnormal{HOT}(\mu_{t},\mu_{v}):=\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}}(\bar{\mu}_{t},\bar{\mu}_{v})$.